In a test your strength booth at a fair you strike a pad with a hammer. This projects a weight up a vertical slide. A bell rings if you have hit the pad hard enough . When the pad is struck with just sufficient force to ring the bell the height h in metres of the weight r seconds after the pad is struck is given h = 8t-4.9r^2.
a) How high is the weight 0.5 seconds after the pad is struck?
b) How high is the bell?
Can you please help me out? Thanks so much in advance:)
Can you also please show the steps it would really help me understand:) Thank-you

Answer by stanbon(75887)

When the pad is struck with just sufficient force to ring the bell the height h in metres of the weight r seconds after the pad is struck is given h = 8t-4.9r^2.
a) How high is the weight 0.5 seconds after the pad is struck?
h(0.5) = 8*0.5 - 4.9(0.5)^2 = 2.775 meters.
---------
b) How high is the bell?
The bell could be 3 miles high. But the weight will reach its
peak when t = -b/(2a) = 8/(2*4.9) = 0.8163 seconds
When t = 0.8163 the height is 8*4 - 4.9*0.8163^2 = 28.73 ft
======================
Cheers,
Stan H.
=====================